The sum of angles of a triangle is 180 degrees. The first angle is 12 degrees more than half the second angle. The third angle is 19 degrees less than 4 times the second angle. What is the measures of all three angles?
Answer by princess08(3)

Let x be the second angle.
Let 12+(x/2) be the first angle.
Let (4*x)-19 be the third angle.
The sum of the angles of a triangle is 180 so,
[12+(x/2)]+x+[(4*x)-19]=180.
12+(x/2)+x+4x-19=180
(x/2)+5x=187
(x/2)+(10x/2)=187
11x=187*2
11x=374
x=34.
12+(x/2)=29.
(4*x)-19=117.
Thus, the angles are 29, 34, and 117.